OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 3x – 2y = 7 9x – 6y = 3 What is the solution to the system? Answer Infinite Solutions (3, 2) No Solutions (3, –2)

hero (hero):

If you want to solve it using graphing technology: y = (7 - 3x)/-2 y = (3-9x)/-6 plot that on a calculator and see what the intersection point is

OpenStudy (anonymous):

i cant... i just kinda need the answer o.o

OpenStudy (mertsj):

The two lines have the same slope and are therefore parallel. They do not intersect so there are no solutions.

hero (hero):

:/ What a shame